Car A leaves sacramento at noon traveling at 60 mph on road 560 miles long to Los Angeles. Car B leaves Los Angeles at 2 pm traveling at constant speed along the smae road to Sacramento. They meet at 6 pm. What was the speed of car B?
Answer by Paul(988)

Since they meet together at same distacne; the distance is 560miles.
Let the speed of the car B be x
Together x+60=560. Now the times for car a is 6h total and the time for car b is 4 hours total.
Car a = 6(60)
car b = 4(x)
together = 560 miles
4(x)+6(60)=560
4x+360=560
4x=200
x=50

Hence, the speed for car B is 50mph.
Paul.